Lorge Corporation has collected the following information after its first year of sales. Sales were $2,500,000 on 100,000 units; selling expenses $250,000 (40% variable and 60% fixed); direct materials $1,360,600; direct labor $260,000; administrative expenses $270,000 (20% variable and 80% fixed); and manufacturing overhead $322,000 (70% variable and 30% fixed). Top management has asked you to do a CVP analysis so that it can make plans for the coming year. It has projected that unit sales will increase by 10% next year. Compute (1) the contribution margin for the current year and the projected year, and (2) the fixed costs for the current year. (Assume that fixed costs will remain the same in the projected year.)

(1) Whats'ts the Contribution margin for current year

What's the Contribution margin for projected year

(2) What' s the ixed costs for current year

Compute the break-even point in units and sales dollars for the first year. (Round contribution margin ratio to 2 decimal places e.g. 0.15 and final answers to 0 decimal places, e.g. 2,510.) Break-even point units Break-even point $
Answer #1
Variable Fixed
Selling Expenses $100,000 $150,000
Direct materials $1,360,600
Direct labor $260,000
Administrative expenses 54,000 216,000
Manufacturing overhead 225400 96600
Total $2,000,000 $462,600

1)

Contribution margin for the current year = $2,500,000 - $2,000,000 = $500,000

Contribution margin for the projected year = ($2,500,000*1.10) -( $2,000,000 * 1.10) = $550,000

2)

Fised cost for the current year and projected year will be same = $462,600

3)

Break even point

Sales = $2,500,000

Unit sales = 100,000

$2,500,000 / 100,000 = $25 per unit sales price

$500,000 / 100,000 = $5 contribution margin per unit

BEP In units = Fixed cost / Contribution per unit

= $462,600 / $5 = 92,520 units

BEP in sales price = BEP in units * sales per unit

= 92,520 * $25 = $2,313,000